OpenStudy (anonymous):

A line segment has one endpoint at the coordinate (-3, 8) and a midpoint at the coordinate (4, 4). Find the other endpoint.

OpenStudy (anonymous):

midpoint = \(\huge \frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\)

OpenStudy (anonymous):

So firstly solve for \(x_2\) here..

OpenStudy (anonymous):

i tried that already it didnt work for me

OpenStudy (anonymous):

\[-3 = \frac{4 + x_2}{2}\] find \(x_2\).. show your steps..

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\(x_2 , y_2\) are the other coordinates that we are to find..

OpenStudy (anonymous):